5. Terri and Mark live 63 miles apart. Both left their homes at 7am by bicycle, riding towards each other. They met at 10am. If Mark’s average speed was three-fourths of Terri’s, how fast did each cycle?
Answer by richwmiller(17219)

m=3t/4
3t+3m=63
3t+3(3t/4)=63
t+3t/4=21
4t+3t=84
7t=84
t=12
m=3t/4
m=9